OpenStudy (donnie1999):

The ratio of men to woman in a certain factory is 3 to 4. There are 228 men. How many workers are there?

OpenStudy (donnie1999):

@crissyyyxoxo

OpenStudy (anonymous):

the ratio means that if there are 3 men there are 4 women . and if there are 6 men there will be 8 women. so if you take that as a fraction you will get men's portion as 3/7 and you can calculate the rest

OpenStudy (donnie1999):

im clueless @Harindu

OpenStudy (anonymous):

3 : 4 228 : ? 228/3 = 76 76 x 4 = 304

OpenStudy (donnie1999):

incorrect :(

OpenStudy (anonymous):

\(\frac{3}{4} = \frac{228}{x} \)

OpenStudy (anonymous):

\(\frac{3}{4} \frac{228}{x} \) Solve for x

OpenStudy (anonymous):

To solve this problem, you write it out as 3/4 = 228/x and then you cross multiply to get 3x=912. Then you solve for x by dividing both sides by three. This will give you the answer that x = 304. So, if there are 228 men, there are 304 women.

OpenStudy (anonymous):

i could of sworn it was right, sorry.